integral one divided by 1+cosx
limits are 0 to pie/4
 

Grade:12th pass

1 Answers

Arijit
6 years ago
1/1+cosx=1/(1+2cos2x/2-1)=1/2sec2x/2 
let tan(x/2)=z
take diff at both sides
½ sec2x/2dx=dz
now∫0(pi/4)dz=[z]0(pi/4)=[tan(x/2)]0(pi/4)=tan(pi/8)